Here is today's summary of NHC computer operations: <BR>--- Matt S.<BR><BR>----------<BR><BR><b>TAFB:</b><p>No new issues.</p><b>HSU:</b><p>1. NCEP model guidance (GFSI, HWFI, HMNI) were not available at 00z.  Monica fixed a path in a script that was modified yesterday and then the models were available before the 03z advisories had to be finished.<BR><BR>2. Did not receive SHIPS/LGEM after submitting guidance for Maria, Lee, and AL99. Also, the SDM reported that they did not receive the inputs for the HWRF/HMON submissions.  Monica investigated and determined that the compute file timestamp needed to be updated and the file resubmitted to fix both problems.<BR><BR>3. All N-AWIPS and ATCF workstations except sawfish and pac-atcf  were rebooted yesterday.</p><b>Other items:</b><p>1. Doug Gaer started adding a number of NHC systems to NCO's Big Brother monitoring<BR></p><b>NHC RFCs:</b><p>1. Monica Bozeman completed RFC 252...Upgrade ATCF FTP scripts to use rsync instead of scp.<BR><BR>2. Doug Gaer completed RFC 3684...NHC video card upgrades for HSU satellite display workstations.<BR></p><BR>